# Towards Open-Vocabulary Audio-Visual Event Localization

### v1

@@ Appendix D Class-wise Performance of the Proposed Two Baselines
In Table 2 of our main paper, we present the overall performance of the proposed training-free and fine-tuning baselines on the test set. Here, we further report their performance on each individual event class. As shown in Fig. 4, the fine-tuning baseline outperforms the training-free baseline in most event classes (approximately 56 out of 67) across all evaluation metrics. This highlights the benefits of additional fine-tuning on training data. Moreover, we observe that some event classes, such as slot machine and chicken crowing, remain challenging for prediction, suggesting avenues for further improvement in future work.
@@
Appendix E More Details on Prompts for adapting Video-LLaMA2 to our OV-AVEL task
@@
In Table 9 of our main paper, we compare the training-free baseline with an advanced audio-visual LLM, namely Video-LLaMA2 [9]. Video-LLaMA2 can process video frames and, more importantly, it can handle general audio signals that are not limited to human speech, unlike other audio-visual LLMs [34]. This makes it particularly suitable for the studied OV-AVEL task. Here, we provide more details on the prompts for adapting Video-LLaMA2 for the OV-AVEL task. Specifically, we tried several prompts and found the following prompt to be the most robust and effective for making predictions: “Instruction: For the given 10-second video, divide it into 10 one-second segments. For each segment, if its audio and visual streams describe the same event, assign the label “x” as “1”; otherwise, label this segment as “0”. User request: After processing all 10 video segments, you will obtain a list with 10 elements, each element being either “1” or “0” according to the above Instruction. Finally, return the most relevant event category of the video from the candidate category list: [“airplane flyby”, “ambulance siren”, “arc welding”, “baby laughter”, “basketball bounce”, “bird chirping”, “bowling impact”, “cat purring”, “cattle mooing”, “chainsawing trees”, “chicken crowing”, …(notebly, all event category texts should be listed; here, we omit the remaining ones for simplicity )]. The output format should be: “ave:” A python list [x, x, x, x, x, x, x, x, x, x] (replace “x” with “1” or “0” according to the prediction); Insert a line break. “class:” the most highly relevant class from the given category list (no punctuation needed at the end).” Readers may directly test this prompt on the official demo website using Hugging Face platform provided by authors of Video-LLaMA2 [9]: https://huggingface.co/spaces/lixin4ever/VideoLLaMA2. In this way, we can obtain the audio-visual event predictions of each test video and compare its performance with the proposed training-free baseline, as reported in Table 9 of our main paper. Additionally, we display some qualitative results in Fig. 5 and Fig. 6 and provide more discussions in Sec. F.
Appendix F Qualitative Results
@@
We finally present some intuitive video examples for OV-AVEL, as shown in Fig. 5 and Fig. 6. Specifically, we visualize the predictions generated by Video-LLaMA2 [9], along with the proposed training-free and fine-tuning baselines. As shown in the figures, the proposed fine-tuning baseline generally yields more accurate temporal localization results for both seen and unseen events/videos. For instance, in the three examples shown in Fig. 5, Video-LLaMA2 tends to predict most video segments as background, indicating its limitation in accurately perceiving the audio-visual correspondence at a fine-grained temporal-level. Although the training-free baseline performs better than Video-LLaMA2, the predictions for some video segments remain unsatisfactory. In contrast, the fine-tuning baseline performs better in localizing temporal segments containing audio-visual events and classifying the event categories. Similar phenomena can also be observed from Fig. 6. These qualitative results, along with the quantitative results presented in our main paper, suggest the effectiveness and superiority of the proposed fine-tuning baseline.

### v2

@@ Appendix F Class-wise Performance of the Proposed Two Baselines
In Table 2 of our main paper, we present the overall performance of the proposed training-free and fine-tuning baselines on the test set. Here, we further report their performance on each individual event class. As shown in Fig. A4, the fine-tuning baseline outperforms the training-free baseline in most event classes (approximately 56 out of 67) across all evaluation metrics. This highlights the benefits of additional fine-tuning on training data. Moreover, we observe that some event classes, such as slot machine and chicken crowing, remain challenging for prediction, suggesting avenues for further improvement in future work.
@@
Appendix G More Details on Prompts for adapting Video-LLaMA2 to our OV-AVEL task
@@
In Table 9 of our main paper, we compare the training-free baseline with an advanced audio-visual LLM, namely Video-LLaMA2 [7]. Video-LLaMA2 can process video frames and, more importantly, it can handle general audio signals that are not limited to human speech, unlike other audio-visual LLMs [34]. This makes it particularly suitable for the studied OV-AVEL task. Here, we provide more details on the prompts for adapting Video-LLaMA2 for the OV-AVEL task. Specifically, we tried several prompts and found the following prompt to be the most robust and effective for making predictions: “Instruction: For the given 10-second video, divide it into 10 one-second segments. For each segment, if its audio and visual streams describe the same event, assign the label “x" as “1"; otherwise, label this segment as “0". User request: After processing all 10 video segments, you will obtain a list with 10 elements, each element being either “1" or “0" according to the above Instruction. Finally, return the most relevant event category of the video from the candidate category list: [“airplane flyby", “ambulance siren", “arc welding", “baby laughter", “basketball bounce", “bird chirping", “bowling impact", “cat purring", “cattle mooing", “chainsawing trees", “chicken crowing", …(notebly, all event category texts should be listed; here, we omit the remaining ones for simplicity )]. The output format should be: “ave:" A python list [x, x, x, x, x, x, x, x, x, x] (replace “x" with “1" or “0" according to the prediction); Insert a line break. “class:" the most highly relevant class from the given category list (no punctuation needed at the end).” Readers may directly test this prompt on the official demo website using Hugging Face platform provided by authors of Video-LLaMA2 [7]: https://huggingface.co/spaces/lixin4ever/VideoLLaMA2. In this way, we can obtain the audio-visual event predictions of each test video and compare its performance with the proposed training-free baseline, as reported in Table 9 of our main paper. Additionally, we display some qualitative results in Fig. A5 and Fig. A6 and provide more discussions in Sec. H.
Appendix H Qualitative Results
@@
We finally present some intuitive video examples for OV-AVEL, as shown in Fig. A5 and Fig. A6. Specifically, we visualize the predictions generated by Video-LLaMA2 [7], along with the proposed training-free and fine-tuning baselines. As shown in the figures, the proposed fine-tuning baseline generally yields more accurate temporal localization results for both seen and unseen events/videos. For instance, in the three examples shown in Fig. A5, Video-LLaMA2 tends to predict most video segments as background, indicating its limitation in accurately perceiving the audio-visual correspondence at a fine-grained temporal-level. Although the training-free baseline performs better than Video-LLaMA2, the predictions for some video segments remain unsatisfactory. In contrast, the fine-tuning baseline performs better in localizing temporal segments containing audio-visual events and classifying the event categories. Similar phenomena can also be observed from Fig. A6. These qualitative results, along with the quantitative results presented in our main paper, suggest the effectiveness and superiority of the proposed fine-tuning baseline.